OpenStudy (anonymous):

If sin A , cos A , tan A are in Geometric Progression . then \[\cot^6 a-\cot^2a\]= ??

OpenStudy (anonymous):

\[\cos^2A=\sin A \tan A=\frac{\sin ^2A}{\cos A}\] \[\cot^2A=\frac{1}{\cos A}\] \[\cot^6A-\cot^2A=\frac{1}{\cos^3A} -\frac{1}{ \cos A}\]

OpenStudy (anonymous):

simplify it to get an integer value!

OpenStudy (anonymous):

how did u get \[\huge \cot^6a=\frac{ 1 }{ \cos^2a }\]?

OpenStudy (anonymous):

\[\frac{\cos A-\cos ^3A}{\cos^4A}=\frac{\cos A(1-\cos^2A)}{\cos^4A}=\frac{ (\sin^2A)}{\cos^3A}\]

OpenStudy (anonymous):

please answer my query !

OpenStudy (anonymous):

since \[\cot ^2A=\frac{1}{\cos A}\] cube everyside... \[(\cot ^2A)^3=(\frac{1}{\cos A})^3\implies \cot ^6A=\frac{1}{\cos^3 A}\]\]

OpenStudy (shubhamsrg):

simplify 1/cos^3 A - 1/cosA and use cos^2 A = sinA tanA to get your finals answer

OpenStudy (anonymous):

i cant find a number 4 that identity...

OpenStudy (shubhamsrg):

There exists an integer for that. Because we know the value of sin^2 A / cos^3 A

OpenStudy (anonymous):

im getting 1

OpenStudy (anonymous):

1/cos^3 A - 1/cosA 1 - cos^2A ------------ cos^3A Tan^2A -------- cos A cos A ---- (cos^2A = sinA tanA => tan^2a = cosA) cos A 1